OpenStudy (anonymous):

Hi you guys! I need some help finding the discriminate of "y=x^2-13x+42.25".. Would anyone like to help? :)

Parth (parthkohli):

1. Scary pic. 2. Did you mean the "discriminant"?

OpenStudy (rulnick):

discriminant is b^2-4ac a 1 b -13 c 42.25

OpenStudy (anonymous):

Is that all that is to it? Yes, I do mean discriminant lol sorry!! And thank you thank you, my pic is very scary I suppose :p

OpenStudy (anonymous):

So when solving, do I leave out "a" and solve with just b and c by plugging them into the above equation?

Parth (parthkohli):

Yes, that is all to it. @rulnick has very nicely given you the values of a, b and c. Using these values, you find what the expression \(b^2 - 4ac\) evaluates to.

Parth (parthkohli):

In a quadratic \(ax^2 + b x+ c = 0\), \(b^2 - 4ac\) is the discriminant.

OpenStudy (anonymous):

I got -169-169, I don't feel as if this is correct :c Am I right?

Parth (parthkohli):

Well, almost so. But \(b^2 = (-13)^2 = 169\) as the square of a negative is also positive.

OpenStudy (anonymous):

ah, because a negative and negative multiplied together creates a positive. So the CORRECT answer would be 169-169? Thank you for helping btw :D

Parth (parthkohli):

Yes... what is 169 minus 169?

OpenStudy (anonymous):

zero

OpenStudy (anonymous):

So the discriminant is zero?

Parth (parthkohli):

yes

OpenStudy (anonymous):

So, would you mind steering me in the next direction? I need to know which would be the best method of finding the x-intercepts algebraically.. I'm not quite sure how to go about knowing this

Parth (parthkohli):

The x-intercept is when the value of the quadratic becomes zero.

OpenStudy (anonymous):

My homework would like to know though, would factoring, quadratic formula, completing the square, etc. be the best method of solving

OpenStudy (anonymous):

So based off of what I have, How am I sure?

Parth (parthkohli):

For this case, you can use any of the above.

OpenStudy (anonymous):

Oh alright, well that's awesome for my assignment! :) Would any of them be a better choice to go with?

Parth (parthkohli):

Really, I'd prefer none over the other. But one thing that's worth a mention: did you know that a quadratic with discriminant zero is a perfect square?

OpenStudy (anonymous):

No, I didn't happen to know this, but I did just get done working with perfect squares awhile ago in my AlgII course and I'm familiar with them. And thank you, I'll just pick one and solve; I appreciate your time! c:

Parth (parthkohli):

No problem! I may prove it to you using the quadratic formula, if you want...

OpenStudy (anonymous):

I would love to watch! Just more demonstration for me, which is helpful always

Parth (parthkohli):

\[x = \dfrac{-b\pm \sqrt{b^2 - 4ac}}{2a}\]Note how that thing within the square root is the discriminant? That's the reason why it's so important to calculate. Anyway, if the discriminant is zero, then\[x = \dfrac{-b \pm \sqrt{0}}{2a}=\dfrac{-b}{2a}\]So basically, we get only one unique root. But you know that a quadratic with only one unique root is a perfect square...

OpenStudy (anonymous):

Oh my, it makes perfect sense now that I see how it is related and comes together

Parth (parthkohli):

Haha!